OpenStudy (anonymous):

Please help with my algebra hw :) WILL FAN AND MEDAL Find the rate of change for each table and equation. How did you fid the rate of change? graph: x= 2 4 6 y= 10 20 30 Equation: 5x+3y=-2

OpenStudy (anonymous):

x=-2 and y=4

OpenStudy (anonymous):

how did you find the rate of change?

OpenStudy (anonymous):

@Data_LG2 do you mind helping me on this problem ?

OpenStudy (anonymous):

ok so i need to turn this in soon and i need to find the rate of change for both the equation and graph. i understand that for the graph the x increases by 2 and the y increases by 10, but how do i write that out on my paper? then i have to explain to my teacher how i got the answer

OpenStudy (anonymous):

There are two methods: Using the slope formula: The formula for slope is: \[\frac{y _{2} - y_{1}} { x_{2} - x_{1}}\ \] Plug-in two points to find the slope: \[\frac{20-10}{4-2} = \frac{10}{2} = 5\] Using y=mx+b form. (Slope-intercept form) \[5x+3y = -2\] to \[3y = 5x -2\] to \[y = 5/3x -2/3\] The slope is m. (5/3)

OpenStudy (anonymous):

@SpooksDF thank you for breaking it down for me that makes a lot of sense.